OpenStudy (anonymous):

Direct Variation? If x varies directly as y, and x = 7 1/2 when y = 10, find x when y = 4. x=?

OpenStudy (bahrom7893):

y=kx

OpenStudy (bahrom7893):

10=k(15/2) 20=15k k=20/15=4/3

OpenStudy (bahrom7893):

y=(4/3)x 4=(4/3)x x=3

OpenStudy (anonymous):

thank you bah!

OpenStudy (bahrom7893):

np becca